# EARN Fairness: Explaining, Asking, Reviewing and Negotiating Artificial Intelligence Fairness Metrics Among Stakeholders

### v1

@@ Abstract.
Numerous fairness metrics have been proposed and employed by artificial intelligence (AI) experts to quantitatively measure bias and define fairness in AI models. Recognizing the need to accommodate stakeholders’ diverse fairness understandings, efforts are underway to solicit their input. However, conveying AI fairness metrics to stakeholders without AI expertise, capturing their personal preferences, and seeking a collective consensus remain challenging and underexplored. To bridge this gap, we propose a new framework, EARN Fairness, which facilitates collective metric decisions among stakeholders without requiring AI expertise. The framework features an adaptable interactive system and a stakeholder-centered EARN Fairness process to Explain fairness metrics, Ask stakeholders’ personal metric preferences, Review metrics collectively, and Negotiate a consensus on metric selection. To gather empirical results, we applied the framework to a credit rating scenario and conducted a user study involving 18 decision subjects without AI knowledge. We identify their personal metric preferences and their acceptable level of unfairness in individual sessions. Subsequently, we uncovered how they reached metric consensus in team sessions. Our work shows that the EARN Fairness framework enables stakeholders to express personal preferences and reach consensus, providing practical guidance for implementing human-centered AI fairness in high-risk contexts. Through this approach, we aim to harmonize fairness expectations of diverse stakeholders, fostering more equitable and inclusive AI fairness.
@@ 1. Introduction
Ensuring artificial intelligence (AI) fairness is crucial for ethical integrity, legal compliance, and human trust (Landers and Behrend, 2023), particularly in high-stakes scenarios such as healthcare (Chien et al., 2022), finance (Kurshan et al., 2022), and university admission (Marcinkowski et al., 2020). AI fairness experts have made significant efforts to define fairness quantitatively through fairness metrics (Mehrabi et al., 2021; Verma and Rubin, 2018). These metrics help assess the fairness of AI systems, audit their compliance with specific fairness requirements, and guide unfairness mitigation (Hort et al., 2023). Currently, there are over 20 popular fairness metrics, categorized into three primary categories — group fairness, subgroup fairness, and individual fairness — to measure whether AI systems treat groups, subgroups, and individuals fairly, respectively (Verma and Rubin, 2018; Mehrabi et al., 2021; Caton and Haas, 2023).
@@
However, there are three primary issues when applying fairness metrics to assess AI fairness. Firstly, choosing the right metrics becomes problematic due to the lack of clear, unified guidelines for metric selection, especially in the absence of a one-size-fits-all metric and the insufficiency of using a single metric (Mehrabi et al., 2021; Madaio et al., 2022; Castelnovo et al., 2022; Haider et al., 2024). Secondly, AI fairness metric selection, often led by AI experts lacking stakeholder input, risks overly depending on these experts’ perspectives and disregards fairness as a societal value with diverse interpretations among stakeholders (Ashktorab et al., 2023; Verma and Rubin, 2018; Mehrabi et al., 2021; Hort et al., 2023). This could lead to AI models deemed technically ”fair” by AI experts but not accepted in practice because they fail to reflect stakeholders’ views on fairness (Whittaker et al., 2018; Lee et al., 2020). Lastly, even when personal fairness perspectives are collected from stakeholders, previous work has shown that different stakeholders often have diverse fairness goals (Madaio et al., 2022; Cheng et al., 2021; Nakao et al., 2023; Lee et al., 2017; Haider et al., 2024). However, AI fairness design should balance multiple stakeholders’ needs and achieve important collective goals (Zhu et al., 2018; Lee et al., 2019a, b). Hence, a conundrum remains: which fairness metrics should we choose that align with stakeholders’ diverse preferences?
Recent studies (Wong, 2020; Cheng et al., 2021; Zhang et al., 2020; Srivastava et al., 2019; Saxena et al., 2019) have recognized the need to identify stakeholders’ metric preferences, for example, Cheng et al. (Cheng et al., 2021) developed an interface with an interview protocol to elicit each stakeholder’s fairness notions in child maltreatment risk management. However, to address the above conundrum, two main research challenges remain. First, existing studies only explored a narrow range of fairness metrics and were hindered by the difficulty of explaining these metrics to stakeholders lacking AI expertise. Secondly, these works focused on investigating individual-level metric preferences, overlooking differences among stakeholders. Although there is a growing recognition of the importance of negotiating views on fairness and establishing consensus (Engelbutzeder et al., 2023; Lee et al., 2019a, b), there are gaps in our knowledge about when stakeholders’ fairness metric preferences collide, whether consensus can be reached on metric selection, what kind of consensus is reached, and how it is reached.
@@
To address the aforementioned challenges and fill the knowledge gaps, we propose the EARN (Explain, Ask, Review, Negotiate) Fairness framework to uncover collective metric preferences among stakeholders without AI knowledge. Our framework includes two components: an adaptable interactive system, the Fairness Explainer and Explorer (FEE), and a stakeholder-centered EARN Fairness process. The FEE communicates eight metrics – six group and subgroup fairness metrics and two individual fairness metrics – to stakeholders. Stakeholders engage with FEE through the two-phase EARN Fairness process: an individual session to Explain fairness metrics to them and Ask for their personal preferences on metrics and reasons, followed by a team session to collectively Review metric definitions and Negotiate a consensus on fairness metrics.
@@
To ground our investigations in a practical context, we applied the EARN Fairness framework in a credit rating model and recruited 18 decision subjects without AI backgrounds. Through individual sessions, we investigated stakeholders’ personal preferences for fairness metrics and their acceptable levels of unfairness—fairness thresholds in AI systems—which serve as critical points where remedial action becomes necessary (Carolyn Saplicki, 2022). Further, using team sessions (collective workshops), we investigated how they established a collective consensus on fairness metrics among them. Through this user study, we address the following research questions:
@@
  • RQ1: What are stakeholders’ personal preferences on AI fairness metrics, and why?
@@
  • We provide a framework for AI practitioners and researchers to engage stakeholders in reaching a collective agreement on metric selection, without AI knowledge being required. This framework can address metric selection by shifting from a technical focus to a stakeholder-centered approach, promoting broader acceptance.
  • We offer AI fairness practitioners empirical results on choosing metrics for a credit rating context that reflect decision subjects’ views, including personal metric preferences and team consensus.
@@
  • We put forward design recommendations for user interfaces (UIs) and fairness explanation methods, enhancing the ability of stakeholders with no AI knowledge to explore AI fairness.
This paper is structured as follows. Section 2 reviews related work on AI fairness metrics, on soliciting stakeholders’ metric preferences, and highlights the research gaps. Next, we introduce the EARN Fairness framework in Section 3. Section 4 describes a user study of applying the framework in a credit rating scenario. We present the results of our user study, answering each research question in Section 5. Finally, in Section 6, we discuss limitations, future research, the generalizability of our results as well as wider implications of our framework.
@@ 2.1. AI Fairness Metrics
Group fairness prioritizes equal treatment among groups, usually defined by protected characteristics enshrined by legislation, such as age or gender, or sensitive features based on contexts, e.g., foreign worker status. Group fairness is easily comprehensible to policymakers and the public, and it is also feasible to implement computationally (Srivastava et al., 2019; Binns, 2020). Some prominent group fairness metrics include (Mehrabi et al., 2021; Hort et al., 2023):
@@
Metrics such as these have been integrated into tools to enable AI experts to assess fairness, such as FairSight (Ahn and Lin, 2020), Silva (Yan et al., 2020; Engelbutzeder et al., 2023), FairVis (Cabrera et al., 2019), and FairRankVis (Xie et al., 2022). Additionally, open-source AI fairness toolkits like IBM’s AI Fairness 360 (Bellamy et al., 2019), Google’s What-If (Wexler et al., 2020), Microsoft’s Fairlearn (Weerts et al., 2023), and DALEX (Baniecki et al., 2021) offer comprehensive approaches for analyzing and mitigating unfairness in AI systems, built around these metrics.
@@ 2.2. Soliciting Stakeholders’ Fairness Metric Preferences
AI fairness experts need to select from numerous metrics in a specific application context, yet there is no agreement on the most appropriate metrics in specific contexts (Hort et al., 2023; Dablain et al., 2022). This has motivated numerous studies to solicit stakeholders’ metric preferences, ensuring AI fairness aligns with societal values (Wong, 2020; Cheng et al., 2021; Saxena et al., 2019; Haider et al., 2024; Ilvento, 2020).
However, explaining metrics to stakeholders without any AI background is challenging due to the complex mathematical concepts of these metrics (Saha et al., 2020). Many studies opt to design various forms of user input to indirectly map users’ metric preferences. One popular method is model selection, where stakeholders’ metric preferences are inferred from their chosen models (Srivastava et al., 2019; Haider et al., 2024; Harrison et al., 2020). For example, lay people have been asked to identify which of two models is more discriminatory based on their prediction labels (Srivastava et al., 2019), or to compare three hypothetical ML models that adhere to different fairness metrics to determine the preferred one (Haider et al., 2024). There are also other forms of input designed to indirectly solicit stakeholders’ metric preferences (Zhang et al., 2020; Saxena et al., 2019; Taka et al., 2023). For example, in loan scenarios, researchers collected stakeholders’ fund allocation decisions to match their preferred metrics (Saxena et al., 2019), or allowed lay users to label AI outcomes as ”fair” or ”unfair” and to adjust feature weights used for model training, thereby calculating which metric has been optimized to uncover their preferences (Taka et al., 2023). These indirect solicitation methods reduce people’s cognitive demands, yet mask the underlying metrics, potentially resulting in choices that inadequately map stakeholders’ true fairness perspectives. Moreover, the user input formats and metrics, along with their application context, are often highly interrelated and need careful design and alignment, thereby lacking the flexibility to adapt to different scenarios.
@@
Once personal preferences for measuring fairness have been extracted, they need to be integrated at a larger scale, to manage differences in preferences. For instance, in loan scenarios, data scientists and domain experts have different perspectives on fairness (Nakao et al., 2023), with data scientists preferring group fairness and domain experts focusing mainly on individual fairness. Even within the same stakeholder types, people hold different metric preferences. For example, in child maltreatment risk management scenarios, some prefer equalized odds, while some prefer statistical parity (Cheng et al., 2021).
Hence, balancing diverse fairness needs and seeking consensus among stakeholders is crucial for implementing fair AI successfully. Current research primarily focuses on seeking consensus on AI model building or predictive outcomes through participatory design (Zhang et al., 2023; Lee et al., 2019b) or co-design (Stumpf et al., 2021; Nakao et al., 2022) for AI fairness democratization. For example, in the context of goods division, Lee et al. (Lee et al., 2019a) reached distribution outcome consensus by allowing individual adjustments first, then group discussions for collective decision-making on the allocations. Furthermore, Lee et al. (Lee et al., 2019b) introduced the WeBuildAI Framework to achieve a consensus-based algorithm driven by voting for fair food donation. To the best of our knowledge, the consensus on fairness metric selection among stakeholders has not been explicitly explored. While recent research advocates and explores negotiating different fairness notions among individuals, it has not translated into practical, quantifiable fairness metrics (Engelbutzeder et al., 2023).
@@ 2.3. Research Gaps
In summary, the current research has three clear gaps. First, the indirect methods used to elicit metric preferences might reduce transparency and hinder stakeholders’ understanding of AI fairness. More detailed explanation work, like that conducted by Cheng et al. (Cheng et al., 2021), which directly interprets metrics, is necessary and should be prioritized to make fairness more comprehensible to a wider audience. Second, due to the complexity of metric explanation, current research generally involves a limited number of metrics and primarily focuses on the popular group fairness category, with little investigation into subgroup and individual fairness. We believe these limitations may not accurately reflect users’ true preferences in certain scenarios. Last, current approaches focus on soliciting personal preferences without considering how to address differences between individuals.
Our work addresses these gaps by designing an adaptive interactive system that visually communicates three fairness categories, including eight metrics, to stakeholders without requiring prior AI knowledge. We aim to uncover stakeholders’ personal preferences, including their fairness threshold settings, and explore how to reconcile differences to reach a metric consensus and what consensus can be achieved.
@@ 3.1.1. Individual Session to Explain and Ask
  • Question 1: What are the three most preferred metrics, and in what order? Given that there is no one-size-fits-all metric and selecting only one metric is insufficient (Mehrabi et al., 2021), we ask each stakeholder to choose and rank their top three fairness metrics, producing a Top 3 Metric List. If group fairness metrics are selected, We further ask whether subgroup fairness for addressing biases across multiple features is more important, and allow them to specify which protected features are of most concern.
  • Question 2: What are the fairness thresholds for different fairness categories? Fairness thresholds quantify the boundaries of stakeholders’ tolerance for unfairness, which is key to aligning AI fairness evaluation with their expectations.
  • Question 3: What are the reasons behind their choices? Stakeholders are encouraged to provide and reflect on the reasons behind their top three metric selections and threshold settings. This process aims to reveal their thought process, allowing a better understanding of their preferences.
@@ 3.1.2. Team Session to Review and Negotiate
Our framework then shifts to a collaborative team session, marrying personal metric preferences into a team consensus among stakeholders. This session involves two key tasks: team members first collectively Review fairness metric definitions using FEE to establish a shared understanding, and then Negotiate. The team needs to accomplish two negotiation objectives. The first objective is to reach a consensus on selecting or defining a fairness metric that aligns with all team members’ fairness expectations. The second objective, when group metrics are adopted, is to determine whether to focus on group fairness for a single feature or subgroup fairness for addressing biases across multiple features.
@@ 3.2.2. Fairness Metrics Covered in FEE
We have expanded the range of fairness metrics beyond prior work in both metric quantity and fairness category (Wong, 2020; Cheng et al., 2021; Zhang et al., 2020; Srivastava et al., 2019; Saxena et al., 2019), introducing eight metrics covering group, subgroup, and individual fairness. Offering a broader range of fairness metrics helps meet the varied expectations and needs of stakeholders and also provides AI experts with more diverse insights into metric selection. Further, given that broader stakeholders, especially those without AI backgrounds, might have divergent understandings of fairness compared to AI experts, our fairness metrics include both widely used metrics in AI fairness practice and relatively less popular ones (Mehrabi et al., 2021; Verma and Rubin, 2018; Hort et al., 2023). See Section 2.1 for definitions and Appendix A.1 for how we implemented these fairness metrics. All metric results are presented as percentages.

### v2

# EARN Fairness: Explaining, Asking, Reviewing, and Negotiating Artificial Intelligence Fairness Metrics Among Stakeholders

@@ Abstract.
Numerous fairness metrics have been proposed and employed by artificial intelligence (AI) experts to quantitatively measure bias and define fairness in AI models. Recognizing the need to accommodate stakeholders’ diverse fairness understandings, efforts are underway to solicit their input. However, conveying AI fairness metrics to stakeholders without AI expertise, capturing their personal preferences, and seeking a collective consensus remain challenging and underexplored. To bridge this gap, we propose a new framework, EARN (Explain, Ask, Review, and Negotiate) Fairness, which facilitates collective metric decisions among stakeholders without requiring AI expertise. The framework features an adaptable interactive system and a stakeholder-centered EARN Fairness process to Explain fairness metrics, Ask stakeholders’ personal metric preferences, Review metrics collectively, and Negotiate a consensus on metric selection. To gather empirical results, we applied the framework to a credit rating scenario and conducted a user study involving 18 decision subjects without AI knowledge. We elicited their personal metric preferences and subsequently we studied how they reached metric consensus in team sessions. Our work shows that the EARN Fairness framework supports stakeholders to express and negotiate fairness preferences, and we provide practical guidance for implementing human-centered AI fairness in high-risk contexts. Through this approach, we aim to reach consensus of fairness perspectives, fostering more equitable and inclusive AI fairness.
@@ 1. Introduction
Ensuring artificial intelligence (AI) fairness is crucial for ethical integrity, legal compliance, and human trust (Landers and Behrend, 2023), particularly in high-stakes scenarios such as healthcare (Chien et al., 2022), finance (Kurshan et al., 2022), and university admission (Marcinkowski et al., 2020). AI fairness experts have made significant efforts to define fairness quantitatively through fairness metrics (Mehrabi et al., 2021; Verma and Rubin, 2018). These metrics help assess the fairness of AI systems, audit their compliance with specific fairness requirements, and guide unfairness mitigation (Hort et al., 2024). Currently, there are over 20 popular fairness metrics, categorized into three primary categories — group fairness, subgroup fairness, and individual fairness — to measure whether AI systems treat groups, subgroups, and individuals fairly, respectively (Verma and Rubin, 2018; Mehrabi et al., 2021; Caton and Haas, 2023).
@@
However, there are three primary issues when applying fairness metrics to assess AI fairness. First, choosing the right metrics becomes problematic due to the lack of clear, unified guidelines for metric selection, especially in the absence of a one-size-fits-all metric and the insufficiency of using a single metric (Mehrabi et al., 2021; Madaio et al., 2022; Castelnovo et al., 2022; Haider et al., 2024). Second, AI fairness metric selection, often led by AI experts lacking stakeholder input, risks overly depending on these experts’ perspectives and disregards fairness as a societal value with diverse interpretations among stakeholders (Ashktorab et al., 2023; Verma and Rubin, 2018; Mehrabi et al., 2021; Hort et al., 2024). This could lead to AI models deemed technically ”fair” by AI experts but not accepted in practice because they fail to reflect stakeholders’ views on fairness (Whittaker et al., 2018; Lee et al., 2020). Previous works (Wong, 2020; Cheng et al., 2021; Zhang et al., 2020; Srivastava et al., 2019; Saxena et al., 2019) have begun to identify stakeholders’ metric preferences, for example, Cheng et al. (Cheng et al., 2021) developed an interface with an interview protocol to elicit each stakeholder’s fairness notions in child maltreatment risk management. However, the range of solicited metrics remains limited, and more work is needed to make fairness metrics more understandable to stakeholders without AI expertise, ensuring fairness is accessible to a broader audience. Last, even when personal fairness perspectives are collected from stakeholders, previous work has shown that different stakeholders often have different fairness goals (Madaio et al., 2022; Cheng et al., 2021; Nakao et al., 2023; Lee et al., 2017; Haider et al., 2024). Although there is a growing recognition of the importance of negotiating views on fairness and establishing consensus (Engelbutzeder et al., 2023; Lee et al., 2019a, b), there are knowledge gaps about when stakeholders’ fairness metric preferences collide, whether consensus can be reached on metric selection, what kind of consensus is reached, and how it is reached. Given these issues, it becomes crucial to not only capture the diverse personal preferences of stakeholders but also investigate fair approaches for negotiating differences and reaching a consensus between multiple stakeholders, ensuring AI fairness aligns with societal values.
@@
To address the aforementioned issues and fill the knowledge gaps, we propose the EARN (Explain, Ask, Review, Negotiate) Fairness framework to elicit personal preferences and support collective negotiation. Our framework includes two components: an adaptable interactive system, the Fairness Explainer and Explorer (FEE), which explains six group and subgroup fairness metrics and two individual fairness metrics to stakeholders, and a stakeholder-centered EARN Fairness process. Stakeholders engage with FEE through the two-phase EARN Fairness process: an individual session to Explain fairness metrics to them and Ask for their personal preferences on metrics and reasons, followed by a team session to collectively Review metric definitions and Negotiate a consensus on fairness metrics.
@@
To ground our investigations in a practical context, we applied the EARN Fairness framework in a credit rating model and recruited 18 decision subjects without AI backgrounds. Through individual sessions, we elicited stakeholders’ personal preferences for fairness metrics. Further, using team sessions, we investigated how they established a collective consensus on fairness metrics among them. Through this user study, we address the following research questions:
@@
  • RQ1: What are stakeholders’ personal preferences on AI fairness metrics that require negotiation?
@@
  • We provide a framework for AI practitioners and researchers to engage stakeholders in reaching a collective agreement on metric selection, without requiring AI knowledge.
@@
  • We put forward design recommendations for user interfaces (UIs) to support the ability of stakeholders with no AI knowledge to explore and negotiate AI fairness.
This paper is structured as follows. Section 2 reviews related work on AI fairness metrics, on soliciting stakeholders’ metric preferences, and highlights the research gaps. Next, we introduce the EARN Fairness framework in Section 3. Section 4 describes a user study of applying the framework in a credit rating scenario. We present the results of our user study, answering each research question in Section 5. Finally, in Section 6, we discuss limitations, future research, as well as wider implications of our research.
@@ 2.1. AI Fairness Metrics
Group fairness prioritizes equal treatment among groups, usually defined by protected characteristics enshrined by legislation, such as age or gender, or sensitive features based on contexts, e.g., foreign worker status. Group fairness is easily comprehensible to policymakers and the public, and it is also feasible to implement computationally (Srivastava et al., 2019; Binns, 2020). Some prominent group fairness metrics include (Mehrabi et al., 2021; Hort et al., 2024):
@@
Metrics such as these have been integrated into tools to enable AI experts to assess fairness, such as FairSight (Ahn and Lin, 2020), Silva (Yan et al., 2020), FairVis (Cabrera et al., 2019), and FairRankVis (Xie et al., 2022). Additionally, open-source AI fairness toolkits like IBM’s AI Fairness 360 (Bellamy et al., 2019), Google’s What-If (Wexler et al., 2020), Microsoft’s Fairlearn (Weerts et al., 2023), and DALEX (Baniecki et al., 2021) offer comprehensive approaches for analyzing and mitigating unfairness in AI systems, built around these metrics.
@@ 2.2. Soliciting Stakeholders’ Fairness Metric Preferences
AI fairness experts need to select from numerous metrics in a specific application context, yet there is no agreement on the most appropriate metrics in specific contexts (Hort et al., 2024; Dablain et al., 2022). This has motivated numerous studies to solicit stakeholders’ personal metric preferences (Wong, 2020; Cheng et al., 2021; Saxena et al., 2019; Haider et al., 2024; Ilvento, 2020)
Explaining metrics to stakeholders without any AI background is challenging due to the complex mathematical concepts of these metrics (Saha et al., 2020). Many studies opt to design various forms of user input to indirectly map users’ metric preferences. One popular method is model selection, where stakeholders’ metric preferences are inferred from their chosen models (Srivastava et al., 2019; Haider et al., 2024; Harrison et al., 2020). For example, lay people have been asked to identify which of two models is more discriminatory based on their prediction labels (Srivastava et al., 2019), or to compare three hypothetical ML models that adhere to different fairness metrics to determine the preferred one (Haider et al., 2024). There are also other forms of input designed to indirectly solicit stakeholders’ metric preferences (Zhang et al., 2020; Saxena et al., 2019; Stumpf et al., 2024). For example, in loan scenarios, researchers collected stakeholders’ fund allocation decisions to match their preferred metrics (Saxena et al., 2019), or allowed lay users to label AI outcomes as ”fair” or ”unfair” and to adjust feature weights used for model training, thereby calculating which metric has been optimized to uncover their preferences (Stumpf et al., 2024). These indirect solicitation methods reduce people’s cognitive demands, yet mask the underlying metrics, potentially resulting in choices that inadequately map stakeholders’ true fairness perspectives. Moreover, the user input formats and metrics, along with their application context, are often highly interrelated and need careful design and alignment, thereby lacking the flexibility to adapt to different scenarios.
@@
Once personal preferences for measuring fairness have been elicited, there is still a need to manage the differences in personal preferences. For instance, in loan scenarios, data scientists and domain experts have different perspectives on fairness (Nakao et al., 2023), with data scientists preferring group fairness and domain experts focusing mainly on individual fairness. Even within the same stakeholder types, people hold different metric preferences, for example, in child maltreatment risk management scenarios, some prefer equalized odds, while some prefer statistical parity (Cheng et al., 2021). Hence, balancing diverse fairness needs and seeking consensus among stakeholders is crucial for implementing fair AI successfully. Current research primarily focuses on seeking consensus on AI model building or predictive outcomes through participatory design (Zhang et al., 2023; Lee et al., 2019b) or co-design (Stumpf et al., 2021; Nakao et al., 2022) for AI fairness democratization. For example, in the context of goods division, Lee et al. (Lee et al., 2019a) reached distribution outcome consensus by allowing individual adjustments first, then group discussions for collective decision-making on the allocations. Furthermore, Lee et al. (Lee et al., 2019b) introduced the ”WeBuildAI” Framework to achieve a consensus-based algorithm driven by voting for fair food donation. While recent research advocates and explores negotiating different fairness notions among individuals, it has not translated into practical, quantifiable fairness metrics (Engelbutzeder et al., 2023). To the best of our knowledge, stakeholders negotiating the consensus on fairness metric selection has not been explicitly explored.
@@ 2.3. Research Gaps
In summary, the current research has two clear gaps. First, indirect methods used to elicit metric preferences might reduce transparency and hinder stakeholders’ understanding of AI fairness. More detailed explanation work, like that conducted by Cheng et al. (Cheng et al., 2021), which directly explains metrics, is necessary and should be prioritized to make fairness more comprehensible to a wider audience. Second, there is a lack of research on how to address differences and achieve consensus between individuals once personal preferences have been elicited.
Our work addresses these gaps by designing an adaptive interactive system that visually communicates three fairness categories, including eight metrics, to stakeholders without requiring prior AI knowledge. Through a novel process in which this interactive system is used, we aim to elicit stakeholders’ personal preferences and then help to reconcile differences to reach consensus on metrics to apply.
@@ 3.1.1. Individual Session to Explain and Ask
  • Question 1: What are the three most preferred metrics, and in what order?
  • Question 2: What are the fairness thresholds for different fairness categories?
  • Question 3: What are the reasons behind their choices?
These questions are used to elicit personal preferences and reasons for these choices to support later negotiation.
@@ 3.1.2. Team Session to Review and Negotiate
Our framework then shifts to a collaborative team session, where personal metric preferences are discussed and negotiated into a team consensus among stakeholders. This session involves two key tasks: team members first collectively Review fairness metric definitions using FEE to establish a shared understanding, and then Negotiate, where they can freely utilize FEE to compare individual preferences and highlight any potential areas of disagreement. The team needs to accomplish selecting or defining a fairness metric that aligns with or reconciles all team members’ fairness preferences.
@@ 3.2.2. Fairness Metrics Covered in FEE
We have expanded the range of fairness metrics beyond prior work in both metric quantity and fairness category (Wong, 2020; Cheng et al., 2021; Zhang et al., 2020; Srivastava et al., 2019; Saxena et al., 2019), introducing eight metrics covering group, subgroup, and individual fairness. Offering a broader range of fairness metrics helps meet the varied expectations and needs of stakeholders and also provides AI experts with more diverse insights into metric selection. Further, given that broader stakeholders, especially those without AI backgrounds, might have divergent understandings of fairness compared to AI experts, our fairness metrics include both widely used metrics in AI fairness practice and relatively less popular ones (Mehrabi et al., 2021; Verma and Rubin, 2018; Hort et al., 2024). See Section 2.1 for definitions and Appendix A.1 for how we implemented these fairness metrics. All metric results are presented as percentages.